Suppose a soccer player boots the ball from a distance 33.5 m toward the goal. find the initial speed of the ball if it just passes over the goal,2.4 m above the ground, given the initial direction of 40 degrees above the horizontal.
Answer #1

Gravitational acceleration = g = -9.81 m/s2

Initial velocity of the ball = V0

Direction of initial velocity = \theta = 40o

Initial horizontal velocity of the ball = Vx0 = V0Cos\theta

Initial vertical velocity of the ball = Vy0 = V0Sin\theta

Horizontal distance of the goal from the initial point = R = 33.5 m

Height of the ball when it crosses the goal = 2.4 m

Time taken by the ball to reach the goal = T

There is no horizontal force acting on the ball therefore the horizontal acceleration is zero.

R = Vx0T

R = (V0Cos\theta)T

T= VoCost

H = Vy0T + gT2/2

H = (VoSino) (1.COsp) + (vCost)?

H = RT and + IR? 2C 24

2.4 = (33.5)Tan(40) +- (-9.81) (33.5) 2V Cos2 (40)

2.4 = 28.11 - 9380.385 V

9380.385 V2 -= 25.71

V02 = 364.853

V0 = 19.1 m/s

Initial speed of the ball = 19.1 m/s